Model with a quadratic equation. Then solve.
A rectangular flower garden is twice as long as it is wide. It has a brick walkway 2 ft wide around it. The area of the flower arden and walkway is 336 ft ^2. find the dimensions of the flower garden.
Please help!

Answer by scott8148(6628)

let x=width , so 2x=length
2 ft walkway on BOTH ends of length and width

(x + 4)(2x + 4) = 336

FOILing ___ 2x^2 + 12x + 16 = 336 ___ x^2 + 6x - 160 = 0

factoring ___ (x + 16)(x - 10) = 0

x = -16 ___ and ___ x = 10 ___ negative value not realistic